NaCl molecule is bound due to electric force between the Sodium and chlorine Ion when one electron of sodium is transferred to chlorine. Taking separation between the ions to be 2.75 * 10^-8 cm, find the force of attraction between them.

Grade:11

2 Answers

Shwetabh
6 years ago
r=2.75×10^-8×10^-2 So,r=2.75×10^-10F=1/4piepsnot×q1q2/r^2F=9×10^-9×1.6×10^-19×1.6×10-19/(2.75×10^-10)^2F=9×10^9×1.6×1.6×10^-38×10-20/2.75×2.75F=9×1.6×1.6×10^29-38/2.75×2.75F=9×1.6×1.6×10^-9/2.75×2.75F=23.04/7.56×10^-9So,F=3.04×10-9 NewtonWe have take one one charge of na and cl because it is mentioned in the question that one electron of sodium is transferred to chlorine So, Na=10electron 11 proton Cl=18 electron 17 protonSo,Na and Cl have one charge
Anjali Shukla
5 years ago
Let the given separation between the ions be 2.75*10-8 cm
Net charge on chloride ion, q1 = 1.6*10-19 C
Net charge on sodium ion, q2 = 1.6*10-19 C
F= kq1q2/r2(sq.)
F = 9*10 9 * (1.6*10-19)2 / (2.75*10-8)2
F = 3.05*10-9 N (Answer)